OpenStudy (anonymous):

f(x) = x^3 + 2x + 6 if A)f^ −1(6) B) f −1(3)

OpenStudy (anonymous):

\[f(0)=6\implies f^{-1}(6)=0\]

myininaya (myininaya):

so we have \[\text{ if } f^{-1}(3)=t => f(t)=3\] so this means we need to solve f(t)=3 for t \[t^3+2t+6=3\] \[t^3+2t+6-3=0\] \[t^3+2t+3=0\] We can guess a solution how about t=-1 -1-2+3=0 So yes! t=-3=>f(-3)=3 => \[f^{-1}(3)=-3\]

myininaya (myininaya):

oops t=-1*

myininaya (myininaya):

\[t=-1=>f(-1)=3=>f^{-1}(3)=-1*\]

OpenStudy (anonymous):

perfect thnk you so much ery good explantation as well :)